# Autoregressive Image Generation Guided by Chains of Thought

### v1

@@ Abstract
In the field of autoregressive (AR) image generation, models based on the ‘next-token prediction’ paradigm of LLMs have shown comparable performance to diffusion models by reducing inductive biases. However, directly applying LLMs to complex image generation can struggle with reconstructing the structure and details of the image, impacting the accuracy and stability of generation. Additionally, the ‘next-token prediction’ paradigm in the AR model does not align with the contextual scanning and logical reasoning processes involved in human visual perception, limiting effective image generation. Chain-of-Thought (CoT), as a key reasoning capability of LLMs, utilizes reasoning prompts to guide the model, improving reasoning performance on complex natural language process (NLP) tasks, enhancing accuracy and stability of generation, and helping the model maintain contextual coherence and logical consistency, similar to human reasoning. Inspired by CoT from the field of NLP, we propose autoregressive Image Generation with Thoughtful Reasoning (IGTR) to enhance autoregressive image generation. IGTR adds reasoning prompts without modifying the model structure or raster generation order. Specifically, we design specialized image-related reasoning prompts for AR image generation to simulate the human reasoning process, which enhances contextual reasoning by allowing the model to first perceive overall distribution information before generating the image, and improve generation stability by increasing the inference steps. Compared to the AR method without prompts, our method shows outstanding performance and achieves an approximate improvement of 20%.
@@ 1 Introduction
Figure 1: Illustration of the reasoning process based on LLMs. The grey circles represent reasoning prompts. (a) General AR model reasoning process. (b) CoT utilizes intermediate textual reasoning prompts to perform text reasoning tasks. (c) Image Generation with Thoughtful Reasoning (IGTR) incorporates image-related reasoning prompts to perform image generation tasks.
@@
In the field of natural language processing (NLP), one of the most distinguishing abilities of LLMs is the exceptional reasoning capabilities, with Chain-of-Thought prompting (CoT) [39] serving as a key technique. Compared to the general AR model reasoning process in Fig. 1 (a), CoT (Fig. 1 (b)) leverages intermediate reasoning prompts to guide the model, achieving significant performance improvements in complex tasks such as arithmetic, commonsense reasoning, and symbolic inference. This structured reasoning approach enhances the stability of the reasoning process and reduces errors caused by the complexity of the task. Moreover, CoT enables the model to reason and solve problems more like humans by incorporating intermediate reasoning prompts, allowing the model to maintain contextual coherence and logical consistency throughout the reasoning process.
Despite differing in task formulation from the logical reasoning tasks commonly encountered in NLP-based LLMs, AR image generation can be attributed to a similar organized structure, where each subtask is solved based on the state of the previous one. As a result, autoregressive image generation also has the powerful reasoning capabilities of LLMs. Therefore, inspired by CoT from the field of NLP, to enhance the ability of LLMs to handle complex image generation tasks, we propose autoregressive Image Generation with Thoughtful Reasoning (IGTR) based on the AR model by adding intermediate reasoning prompts for image generation, without modifying any model structure or the autoregressive raster generation order, which is shown in Fig. 1 (c). Our method improves both the accuracy and stability of the generation and strengthens the logical generation capabilities of the images. Specifically, we design specialized image-related reasoning prompts for AR image generation as the intermediate reasoning process, simulating the human reasoning process by allowing the model to first perceive overall distribution information (such as the universal distribution of the image dataset or the overall distribution information of a specific category) before generating the image. The reasoning prompt enhances contextual reasoning of generation, and also improves stability during the generation process by increasing the inference step.
Our contributions are as follows:
  1. 1.
    We propose autoregressive Image Generation with Thoughtful Reasoning (IGTR), where we introduce CoT technology to autoregressive image generation by incorporating intermediate reasoning prompts for image generation. IGTR enhances the capability of large language models in the complex image generation task, without modifying any model structure or the autoregressive raster generation order.
  2. 2.
    We design specialized image-related reasoning prompts for AR image generation as the intermediate reasoning process, allowing the model to first perceive overall distribution information before generating the image, which enhances contextual reasoning of generation, and also improves stability during the generation process by increasing the inference step.
  3. 3.
    Compared to the AR method without prompts, our method shows outstanding performance and achieves an approximate improvement of 20%.
@@ 2 Related Work
2.1 Chain-of-Thought
Large language models (LLMs) exhibit impressive reasoning capabilities due to advanced language generation abilities. Chain-of-Thought (CoT), as a key technique, is specifically designed for multi-step reasoning [39]. Traditional prompting methods [3] have shown limited effectiveness on complex tasks that require reasoning capabilities. Inspired by the concept of using intermediate steps to address reasoning problems [5, 20], CoT simulates the human process of reasoning and problem-solving by incorporating intermediate reasoning prompts, allowing the model to maintain contextual coherence and logical consistency during the reasoning process [39]. CoT has been shown to significantly enhance the performance of LLMs on complex tasks, while also improving the stability of the reasoning process [39].
Chain-of-Thought has been applied to many different tasks, including natural language processing [21, 38, 40, 41] and multi-modal large language models [45, 22, 23]. To the best of our knowledge, in the field of AR image generation, the capabilities of CoT have not yet been explored, which can provide the intermediate reasoning process for complex image generation tasks, to enhance image generation capabilities, strengthen the logical generation of images, and improve the stability of generation.
@@ 2.2 Autoregressive Visual Generation Methods
Visual generation has received significant attention in recent years, especially with the development of deep learning architectures and generative models. One key trend is the use of autoregressive models, which show significant generality and potential due to their strong connection with NLP. However, there is no mature or well-established community, and further efforts are needed to overcome challenges and fully realize their capabilities. Existing autoregressive generation methods are commonly divided into two approaches: BERT-style mask autoregressive models and GPT-style autoregressive models. Mask Autoregressive methods [4, 43, 18, 19], inspired by BERT-style pre-training [15], generate images by predicting the random masked tokens. Instead, another kind of autoregressive method [8, 42, 26], inspired by GPT [25], predicts the next token in a sequence, which applies the image tokenization [16, 37] to transform images to discrete space. Recently, based on autoregressive methods, LlamaGen [35] adapts large language model architectures like Llama [36], to autoregressively predict image tokens by applying the ‘next-token prediction’ paradigm to visual generation, which achieve decent performance in image generation.
@@ 2.3 Other Visual Generation Methods
In addition to AR visual generation methods, significant efforts have also been made in exploring other forms of visual generation models. Generative Adversarial Networks (GANs) is the earliest approaches, leveraging adversarial training to generate images [10, 2, 13, 30, 14]. Diffusion models, an alternative approach, generates images by gradually refining random noise through a series of learned steps [33, 32, 6, 12, 27, 24]. GANs-based and diffusion-based methods show promising performance because their community are relatively complete. If AR models are to surpass them, further iterations and development of AR models are needed.
@@ 3.2.1 Overview
Our method still adapts the image tokenizer architecture and ‘next-token prediction’ prediction form of the classic AR model. On this basis, we introduce a set of reasoning prompt $\tilde{S}=(s_{1},s_{2},...,s_{k})$ in the training and inference process, just like the CoT technique in the language model, and modify the model $\theta$ training target to maximize the log-likelihood function 2. The way model inference is change to $\prod^{T}_{t=1}P(x_{t}|x_{<t},\tilde{S},c)$.
@@ 3.2.2 Theoretical Analysis
The introduced resoning prompt is $\tilde{S}=(s_{1},s_{2},...,s_{k})$, the generated token sqeuence is $\tilde{X}=(x_{1},x_{2},...,x_{t})$, the condition is $c$, and the model parameter is $\theta$.
Figure 2: The exhibition of different state with k-length reasoning prompt.
@@
We refer to the derivation of CoT techniques for LLMs in the context of general dynamic programming problems by Feng et al. [9], and analyse the impact of introducing CoT in our model while training. As shown in Figure 2, one sequence shows the different state $i$ with unidirectional transitions between states constrained by the time step $i$. Since there is a target sequence present, for each token prediction during model training, there exists an optimal solution token (i.e., the target token at the corresponding position). Thus, we can express the state transition function as follows:
@@
$p_{j}$ represents the probability of the model predicting the optimal state $j$ token at time step $j$. $T(i,j)$ denotes the probability of transitioning from the optimal state $i$ to the optimal state $j$, where the optimal state is defined as a state in which all tokens are the optimal tokens for current position. Thus, we can formulate $T_{AR}(1,t)$ for the classic AR model and $T_{IGTR}(1,t)$ for our method as follows:
@@
It is easy to infer that, under the ideal condition with the same target, both $T_{AR}(1,t)$ and $T_{IGTR}(1,t)$ should equal to the expression $\prod^{T}_{t=1}P(x_{t}|x_{<t},c)$, which corresponds to the maximum likelihood estimation (MLE) of the target generated image. However, since the limitation of the model ability, what we can do is to approximates the target distribution through MLE, these tree equations $T_{AR}(1,t)$, $T_{IGTR}(1,t)$ and $\prod^{T}_{t=1}P(x_{t}|x_{<t},c)$ are not exactly equal. In this case, due to the introduction of reasoning prompt, our method, compared to the classic AR method, gains an additional term $T(2,k+1)$, which allows the model have a set of weights that can be dynamically adjusted based on the condition $c$. Therefore, compared to $T_{AR}(1,t)$, $T(2,k+1)$ makes $T_{IGTR}(1,t)$ more easily approximated to the target distribution, enabling a model with the same parameter count to converge faster when trained on the same amount of data.
The final loss of the model is constrained by the cross-entropy between the target sequence and the generated sequence $\tilde{X}$.
Figure 3: Our method first receives Class Token and Reasoning Prompt as input, and then follows the ‘next-token prediction’ to generate images.
@@
We analyse the impact of introducing reasoning prompt on the model’s inference phase from the perspective of information entropy. The information entropy of the classic AR model during inference can be expressed as $H(\tilde{X}|c)$, while the information entropy of our method with reasoning prompt during inference can be expressed as $H(\tilde{X}|\tilde{S},c)$. With the definition of entropy, we can derive Eq 6 as follows:
@@
This inequality suggests that, by incorporating reasoning prompt, compared to classic AR model, our method has less uncertainty during inference, which means the generation process is more stable. This reduction in uncertainty helps effectively minimize inconsistencies in the image details represented by the tokens, leading to higher-quality generated images. The more stable token generation process ensures that the generated sequence better preserves the coherence of the visual structure, resulting in images with fewer artifacts and better alignment with the target distribution.
3.2.3 Reasoning Prompt
As shown in Figure 3, we choose two different type of reasoning prompt $\tilde{S}$ to incorporate into the model’s training and inference process.
@@ 3.2.3 Reasoning Prompt
    We randomly select an image that is different from the training target while it’s still under the same class with the training target. After passing it through the image tokenizer, it is converted into tokens, which are then concatenated with the condition $c$ and input into the model. During inference, we randomly choose an image, tokenize it, and use the resulting tokens as input directly.
@@
These two different reasoning prompts represent two distinct distributions used to guide the model in generating images. The first prompt corresponds to a specific category distribution called Class Reasoning Prompt, aimed at helping the model better understand the distribution corresponding to the given condition, thereby generating images that more clearly represent the category information. The second prompt represents a more universal image distribution called Universal Reasoning Prompt, designed to encourage the model to generate more diverse and varied images.
@@
Non Reasoning Prompt 5.46 193.61 0.83 0.45
Class Reasoning Prompt Class Reasoning Prompt 4.34 226.31 0.84 0.46
Universal Reasoning Prompt 4.36 227.44 0.84 0.46
Universal Reasoning Prompt Class Reasoning Prompt 4.42 217.66 0.85 0.45
Universal Reasoning Prompt 4.39 222.80 0.85 0.46
@@
Table 2: Ablation study for different types of reasoning prompts using for training or inference. $\downarrow$’ or $\uparrow$’ indicate lower or higher values are better. ‘Mixture’ means we combine two different prompts for training and inference,
@@ 4 Experiment
In this section, we describe the implementation details of our approach in Sec. 3.2.3. We then provide ablation studies on important design decisions in Sec. 4.2. Following that, the main results and the visualizations are discussed in Sec. 4.3.
@@ 4.1 Experiment Setting
Following prior work which use a VQ tokenizer to tokenize the input images into discrete tokens, we use the VQGAN reported in LlamaGen [35] with the official weight trained on ImageNet. Our model architecture is largely based on Llama [36], applying pre-normalization using RMSNorm [44], SwiGLU [31] activation function, and 2D rotary positional embeddings [34] at each layer of our model. We use model of different configuration including B ( [^1] M), L (343M), XL (775M), and XXL (1.4B). The detailed architecture configuration and size are shown in Table 1.
@@
We experiment on ImageNet [28] at a resolution of 256$\times$256 based on the class to image task. To accelerate training, we pre-tokenize the entire training dataset before training using the VQGAN tokenizer reported in LlamaGen [35]. Additionally, we improve the diversity of the pre-tokenized dataset by applying the ten-crop transformation [35]. Following the classic evaluation suite provided by [7], we evaluate FID [11] as the main metric and also report IS [29], Precision, and Recall as references metrics. Unless otherwise specified, models are trained for 300 epochs and the default inference setting is top-k=0, top-p=1.0, temperature=1.0, number of reasoning prompt=256. Additionally, the dropout for the class condition embedding in Classifier-Free Guidance (CFG) is set to 1.75 by default. All models are trained using the base learning rate of 0.0001 and the batch size of 256. AdamW optimizer is set with $\beta_{1}=0.9$ and $\beta_{2}=0.95$, weight decay is set to 0.05, and gradient clipping is set at 1.0. A dropout rate of 0.1 is applied to the input token embeddings, attention modules, and the FFN module.

### v3

# Autoregressive Image Generation with Vision Full-view Prompt

@@ Abstract
In autoregressive (AR) image generation, models based on the ‘next-token prediction’ paradigm of LLMs have shown comparable performance to diffusion models by reducing inductive biases. However, directly applying LLMs to complex image generation can struggle with reconstructing the image’s structure and details, impacting the generation’s accuracy and stability. Additionally, the ‘next-token prediction’ paradigm in the AR model does not align with the contextual scanning and logical reasoning processes involved in human visual perception, limiting effective image generation. Prompt engineering, as a key technique for guiding LLMs, leverages specifically designed prompts to improve model performance on complex natural language processing (NLP) tasks, enhancing accuracy and stability of generation while maintaining contextual coherence and logical consistency, similar to human reasoning. Inspired by prompt engineering from the field of NLP, we propose Vision Full-view prompt (VF prompt) to enhance autoregressive image generation. Specifically, we design specialized image-related VF prompts for AR image generation to simulate the process of human image creation. This enhances contextual logic ability by allowing the model to first perceive overall distribution information before generating the image, and improve generation stability by increasing the inference steps. Compared to the AR method without VF prompts, our method shows outstanding performance and achieves an approximate improvement of 20%.
@@ 1 Introduction
Figure 1: Illustration of the prompt engineering for AR model. The grey module represents prompts. (a) Prompt engineering utilizes designed prompts to perform text tasks. (c) Vision Full-view (VF) Prompt incorporates image-related prompts to perform image generation tasks.
@@
In the field of natural language processing (NLP), prompt engineering is a series of techniques that utilize specific designed prompts to enhance the model’s generative ability [30, 5, 40]. As shown in Fig. 1 (a), compared to the general AR model process, prompt engineering leverages specific designed prompts to guide the model, achieving significant performance improvements in complex tasks such as arithmetic, code generation, and knowledge-intensive question answering. On the one hand, prompt engineering provides clearer and more informative guidance, reducing objective function drift during training and enhancing training stability. During inference, it reduces uncertainties caused by input, thereby improving the stability of model predictions. On the other hand, prompt engineering introduces logical guidance and in-context information for input, enabling the model to better simulate human problem-solving approaches. This enhances the coherence of the reasoning process, reduces erroneous reasoning paths, and improves the interpretability and reliability of generated answers.
Despite differing in task formulation from the logical reasoning tasks commonly encountered in NLP-based LLMs, AR image generation can be attributed to a similar organized structure, where each subtask is solved based on the state of the previous one. Therefore, inspired by prompt engineering from the field of NLP, to handle complex image generation tasks, we propose a Vision Full-view (VF) prompt based on the AR model, which guides the model to grasp the overall visual information before progressively genrate local details during generation, as illustrated in Fig. 1 (b). Specifically, we design specialized VF prompt obtained for AR image generation, simulating the human process of creating images by guiding the model to first perceive vision full-view information (such as the universal distribution of the image dataset or representative full-view visual patterns from a specific category) before generating the image. Therefore, the VF prompt enhances the contextual logical capability of the model and improves generation stability by increasing inference steps. Experimental results demonstrate that our method significantly outperforms the baseline AR method without prompts, achieving an approximate performance improvement of 20%.
@@ 2.1 Autoregressive Visual Generation Methods
Visual generation has received significant attention in recent years, especially with the development of deep learning architectures and generative models. One key trend is the use of autoregressive models, which show significant generality and potential due to their strong connection with NLP. However, there is no mature or well-established community, and further efforts are needed to overcome challenges and fully realize their capabilities. Existing autoregressive generation methods are commonly divided into two approaches: BERT-style mask autoregressive models and GPT-style autoregressive models. Mask Autoregressive methods [4, 45, 19, 20], inspired by BERT-style pre-training [15], generate images by predicting the random masked tokens. Instead, another kind of autoregressive method [9, 44, 27], inspired by GPT [25], predicts the next token in a sequence, which applies the image tokenization [16, 39] to transform images to discrete space. Recently, based on autoregressive methods, LlamaGen [37] adapts large language model architectures like Llama [38], to autoregressively predict image tokens by applying the ‘next-token prediction’ paradigm to visual generation, which achieves decent performance in image generation.
@@ 2.2 Other Visual Generation Methods
In addition to AR visual generation methods, significant efforts have also been made in exploring other forms of visual generation models. Generative Adversarial Networks (GANs) are the earliest approaches, leveraging adversarial training to generate images [10, 2, 13, 32, 14]. Diffusion models, an alternative approach, generates images by gradually refining random noise through a series of learned steps [35, 34, 6, 12, 28, 24]. GANs-based and diffusion-based methods show promising performance because their community are relatively complete. If AR models are to surpass them, further iterations and development of AR models are needed.
2.3 Prompt Engineering for LLMs
Prompt engineering, a critical methodology for optimizing interactions with large language models (LLMs), strategically designs structured instructions to guide models toward high-quality outputs. Previous research has extensively explored various prompting strategies. For example, Zero-shot prompting [26] guides the model to complete a task solely through task descriptions, relying on its pre-trained knowledge for reasoning and generation. Few-shot prompting [3] provides a small number of examples in addition to the task description, helping the model learn task patterns and improving its performance on specific tasks. Chain-of-Thought (CoT) prompting [41, 21, 43] enhances performance in complex tasks such as mathematics and logical reasoning by instructing the model to generate a step-by-step reasoning process (e.g., ‘Let’s think step by step’). In-context Learning (ICL) prompting [22, 8, 23] enables LLMs to learn task patterns directly from the provided context during inference by including task examples or relevant information in the input, eliminating the need for model parameter updates. Retrieval-Augmented Generation (RAG) [18, 42] combines information retrieval systems to dynamically fetch external knowledge during inference. This approach allows the model to generate more factually accurate responses and reduces hallucinations.
The reasons why prompt engineering achieves excellent success in the LLM are as follows. First, it refines input instructions to establish a more structured and consistent learning signal, thereby mitigating shifts in the objective function during training and improving overall model stability. Additionally, during inference, it helps reduce variability caused by input fluctuations, thereby improving the stability of model predictions. Second, prompt engineering introduces logical guidance and in-context information directly into the input enabling the model to align more closely with human-like problem-solving strategies, improving the fluency and coherence of generated responses while minimizing logical inconsistencies and hallucinations. Based on these two reasons, we apply prompt engineering to autoregressive (AR) image generation to address issues of training instability and inconsistency with human perception in AR image generation.
@@ 3.2.1 Overview
Our method still adapts the image tokenizer architecture and ‘next-token prediction’ prediction form of the classic AR model. On this basis, we introduce a set of VF prompts $\tilde{S}=(s_{1},s_{2},...,s_{k})$ in the training and inference process, and modify the model $\theta$ training target to maximize the log-likelihood function 2. The way model inference is change to $\prod^{T}_{t=1}P(x_{t}|x_{<t},\tilde{S},c)$.
@@ 3.2.2 Theoretical Analysis
The introduced VF prompt is $\tilde{S}=(s_{1},s_{2},...,s_{k})$, the generated token sqeuence is $\tilde{X}=(x_{1},x_{2},...,x_{t})$, the condition is $c$, and the model parameter is $\theta$.
Figure 2: The exhibition of different states with k-length Full-view prompt.
@@
We refer to the derivation of prompt engineering for LLMs in the context of general dynamic programming problems and analyse the impact of introducing prompt engineering in our model while training. As shown in Figure 2, one sequence shows the different states $i$ with unidirectional transitions between states constrained by the time step $i$. Since there is a target sequence present, for each token prediction during model training, there exists an optimal solution token (i.e., the target token at the corresponding position). Thus, we can express the state transition function as follows:
@@
$p_{j}$ represents the probability of the model predicting the optimal state $j$ token at time step $j$. $T(i,j)$ denotes the probability of transitioning from the optimal state $i$ to the optimal state $j$, where the optimal state is defined as a state in which all tokens are the optimal tokens for the current position. Thus, we can formulate $T_{AR}(1,t)$ for the classic AR model and $T_{VFprompt}(1,t)$ for our method as follows:
@@
It is easy to infer that, under the ideal condition with the same target, both $T_{AR}(1,t)$ and $T_{VF\_prompt}(1,t)$ should equal to the expression $\prod^{T}_{t=1}P(x_{t}|x_{<t},c)$, which corresponds to the maximum likelihood estimation (MLE) of the target generated image. However, since the limitation of the model ability, what we can do is to approximate the target distribution through MLE, these tree equations $T_{AR}(1,t)$, $T_{VF\_prompt}(1,t)$ and $\prod^{T}_{t=1}P(x_{t}|x_{<t},c)$ are not exactly equal. In this case, due to the introduction of VF prompt, our method gains an additional term $T(c,k)$ compared to the classic AR method.
The final loss of the model is constrained by the cross-entropy between the target sequence and the generated sequence $\tilde{X}$. While it’s not involved in the cross-entropy calculation of classic AR models, it‘s included in VF prompt’s cross-entropy calculation, formatted as $\sum\log T(c,k)$. This term, computed by the model, serves as a bias varying with the VF prompt to help the model avoid instability caused by mismatch between training data and inference scenarios. It maintains a consistent optimization direction across different targets, making it faster to converge to an ideal performance.
Figure 3: Our method first receives Class Token and Vision Full-view prompt as input, and then follows the ‘next-token prediction’ to generate images.
@@
We analyse the impact of introducing VF prompt on the model’s inference phase from the perspective of information entropy. The information entropy of the classic AR model during inference can be expressed as $H(\tilde{X}|c)$. In contrast, the information entropy of our method with VF prompt during inference can be expressed as $H(\tilde{X}|\tilde{S},c)$. With the definition of entropy, we can derive Eq 6 as follows:
@@
This inequality suggests that, by incorporating VF prompt, our method has less uncertainty during inference compared to classic AR model, which means the generation process is more stable. This reduction in uncertainty helps effectively minimize inconsistencies in the image details represented by the tokens, leading to higher-quality generated images. The more stable token generation process ensures that the generated sequence better preserves the coherence of the visual structure, resulting in images with fewer artifacts and better alignment with the target distribution.
3.2.3 Vision Full-view Prompt
As shown in Figure 3, we choose two types of vision full-view prompt $\tilde{S}$ to incorporate into the model’s training and inference process.
@@ 3.2.3 Vision Full-view Prompt
    We randomly select an image different from the training target while it’s still under the same class with the target. After passing it through the image tokenizer, it is converted into tokens, which are then concatenated with the condition $c$ and input into the model. During inference, we randomly choose an image, tokenize it, and use the resulting tokens as input directly.
@@
These two different vision full-view prompts represent two distinct representations used to guide the model in generating images. The first prompt corresponds to a specific category representation called Class VF prompt, aimed at helping the model better understand the distribution corresponding to the given condition, thereby generating images that more clearly represent the category information. The second prompt, called Universal VF prompt, represents a more universal image distribution and is designed to encourage the model to generate more diverse and varied images, independent of the condition.
@@
Non VF prompt 5.46 193.61 0.83 0.45
Class VF prompt Class VF prompt 4.34 226.31 0.84 0.46
Universal VF prompt 4.36 227.44 0.84 0.46
Universal VF prompt Class VF prompt 4.42 217.66 0.85 0.45
Universal VF prompt 4.39 222.80 0.85 0.46
@@
Table 2: Ablation study for different types of VF prompts used for training or inference. $\downarrow$’ or $\uparrow$’ indicate lower or higher values are better. ‘Mixture’ means we combine two different prompts for training and inference,
@@ 4 Experiment
In this section, we describe the implementation details of our approach in Sec. 4.1. We then provide ablation studies on important design decisions in Sec. 4.2. Following that, the main results and the visualizations are discussed in Sec. 4.3.
@@ 4.1 Experiment Setting
Following prior work which uses a VQ tokenizer to tokenize the input images into discrete tokens, we use the VQGAN reported in LlamaGen [37] with the official weight trained on ImageNet. Our model architecture is largely based on Llama [38], applying pre-normalization using RMSNorm [46], SwiGLU [33] activation function, and 2D rotary positional embeddings [36] at each layer of our model. We use models of different configurations including B ( [^1] M), L (343M), XL (775M), and XXL (1.4B). The detailed architecture configuration and size are shown in Table 1.
@@
We experiment on ImageNet [29] at a resolution of 256$\times$256 based on the class to image task. To accelerate training, we pre-tokenize the entire training dataset before training using the VQGAN tokenizer reported in LlamaGen [37]. Additionally, we improve the diversity of the pre-tokenized dataset by applying the ten-crop transformation [37]. Following the classic evaluation suite provided by [7], we evaluate FID [11] as the main metric and also report IS [31], Precision, and Recall as references metrics. Unless otherwise specified, models are trained for 300 epochs and the default inference setting is top-k=0, top-p=1.0, temperature=1.0, number of VF prompt=256. Additionally, the dropout for the class condition embedding in Classifier-Free Guidance (CFG) is set to 1.75 by default. All models are trained using the base learning rate of 0.0001 and the batch size of 256. AdamW optimizer is set with $\beta_{1}=0.9$ and $\beta_{2}=0.95$, weight decay is set to 0.05, and gradient clipping is set at 1.0. A dropout rate of 0.1 is applied to the input token embeddings, attention modules, and the FFN module.